OpenStudy (maddss):

What are the vertices of the hyperbola (x-1)^2/16-(y+2)^2/4=1 A) (3, -2) and (-1, -2) B) (1, 2) and (1, -6) C) (-3, -2) and (5, -2) D) (1, 0) and (1, -4)

OpenStudy (anonymous):

try plugging them into the formula and the ones which give u the corresponding y or x values should be the verticies

OpenStudy (maddss):

whats the formula

OpenStudy (mertsj):

What is the center of your hyperbola?

OpenStudy (mertsj):

Are you there maddss?

OpenStudy (maddss):

Thats the whole question.

OpenStudy (mertsj):

Can you tell the center of the hyperbola by looking at its equation?

OpenStudy (maddss):

no, this is the first work sheet for this chapter in my math class..so i hardly know it

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(x-k)^2}{b^2}=1\]

OpenStudy (mertsj):

You want to make your equation match up with that one. The center is (h,k) so when yours matches up, you will know the center of your hyperbola.

OpenStudy (mertsj):

\[\frac{(x-1)^2}{16}-\frac{(y- -2)^2}{4}=1\]

OpenStudy (mertsj):

What is the center?

OpenStudy (maddss):

-1,2

OpenStudy (mertsj):

Close....(-1,-2)

OpenStudy (mertsj):

Now you need the value of a. What is it?

OpenStudy (maddss):

Are you sure its (-1,-2) you put plugged in the problem like \[\frac{ \left( x-1 \right)^{2} }{ 16 }-\frac{ \left( y- -2 \right)^{2} }{ 4 }\] the (--2) that changes into a positive.

OpenStudy (mertsj):

Please notice the equation of the hyperbola I posted whose center is (h,k) Please note that h and k are the numbers that are AFTER the negative sign. That is why I wrote y+2 in the form y--2...to emphasize that the -2 is the number that is AFTER the negative sign and thus it is the y coordinate of the vertex.

OpenStudy (maddss):

ok

OpenStudy (mertsj):

So back to my question...what is the value of a?

OpenStudy (maddss):

16

OpenStudy (mertsj):

That would be a^2 if you look at the equation I posted you see that the number under the x is a^2 not a

OpenStudy (maddss):

You asked for the letter "a" not the denominator. \[a ^{2}\] is different so that would be 265

OpenStudy (mertsj):